OpenStudy (anonymous):

Need help, thanks I have a review kind of thing which shows this question: 1/6 as percent: 1/6 =x/100 x=16 4/6 = 16 2/3 Isn't it 16.6(repeating) percent? How did they get the mixed fraction 16 4/6 ? Thanks! :)

OpenStudy (anonymous):

*16 2/3%

OpenStudy (scienceteen8):

i wish i was good at fractions

OpenStudy (caozeyuan):

yes16.6% is correct

OpenStudy (jdoe0001):

if you use 100/6 = x you'd end up with 16 and 4/6, thus but that's correct

OpenStudy (jdoe0001):

and \(\bf 16.\overline{16}\iff 16\frac{4}{6}\iff 16\frac{2}{3}\)

OpenStudy (mathmale):

Wouldn't 16 followed by the repeating decimal .16 come out to 16 1/6, not 16 4/6?

OpenStudy (jdoe0001):

ohh yeah shoot \(\bf 16.\overline{6}\iff 16\frac{4}{6}\iff 16\frac{2}{3}\)